# Universality of Efimov states in highly mass-imbalanced cold-atom mixtures with van der Waals and dipole interactions

### v2

@@ III Numerical Results
We show the numerical solutions of Eq. (1) in this section. As we are particularly interested in Er-Er-Li three-body systems, we perform the calculations with Er-Li mass ratio: $M/m=27.5855...$ for the bosonic system corresponding to 166Er-6Li, and $M/m=27.7520...$ for the fermions corresponding to 167Er-6Li, respectively. The differential equation (1) is discretized in the log-scaled coordinate space betwen $R_{\mathrm{min}}$ and $R_{\mathrm{max}}$, where $R_{\mathrm{max}}$ is typically taken as $\sim$ 400-1000 $r_{\mathrm{vdw}}$. We find that 4000 grid points and $\ell_{\mathrm{max}}=$10-20 are enough to attain convergence for most cases, while we take up to 30000 grid points and $\ell_{\mathrm{max}}\approx 40$ for some cases of stronger dipole interactions.
@@ III.1 Non-dipolar case $a_{dd}=0$
Figure 2: Binding energy of the Efimov states plotted against (a)(b) the $s$-wave scattering length $a^{(\mathrm{HH})}$ for the bosons, and (c)(d) $p$-wave scattering volume $v^{(\mathrm{HH})}$ for the fermions, obtained by converting $R_{\mathrm{min}}$ in Fig. 1 into $a^{(\mathrm{HH})}$ and $v^{(\mathrm{HH})}$ using Eqs. (8) (13) (14). The left (right) pannels are negative (positive) scattering length/volume sides, which should smoothly connect with each other in the unitary limit $01/a^{(\mathrm{HH})}=$ 0, $01/v_{p}^{(\mathrm{HH})}=$ 0 corresponding to the border between the left and right panels. The solid curves (black) are the analytical results of Eqs. (15) (16), and the dashed curves (green) are the dimer binding energy of the heavy atoms. The vertical dotted lines (black) are the same as those in Fig. 1
@@
First, we show the results without the dipole interaction: in Fig. 1, we show the binding energy of the Efimov states (in the van der Waals energy unit $1/Mr_{\mathrm{vdw}}^{2}$) as the short-range boundary $R_{\mathrm{min}}$ is varied. As $R_{\mathrm{min}}$ gets smaller, the short-range van der Waals attraction increases, leading to the successive appearance of bound states and increase of the binding energy. The numerical data points almost perfectly agree with the analytical curves obtained with the QDT in Eq. (7), particularly at small energy. There are a few points which significantly deviate from the analytical curves. They are ascribed to higher angular momentum states of $\ell=2,4,...$ for the bosons and $\ell=3,5,...$ for the fermions, respectively. The thresholds at which the $\ell=2$ ($\ell=3$) bound states start to appear, denoted by the vertical dotted lines in Fig. 1 (a) (and (b)), agree with most of the stray data points, supporting this interpretation. A few stray points which agree with neither are ascribed to higher angular momentum channels $\ell\geq 4$ ($\ell\geq 5$). This classification of the states has also been confirmed from the shape of the wavefunctions. Because the higher angular momentum states are localized to shorter distances than the $s$-wave ($p$-wave) state owing to the centrifugal barrier, they are more sensitive to $R_{\mathrm{min}}$, resulting in the steeper binding energy curves. The binding energy curves of different partial waves do not render an avoided crossing but simply cross because the angular momentum is a good quantum number in the absence of the dipole interaction.
@@
In Fig. 2, we show the binding energies plotted against the $s$-wave scattering length $a^{(\mathrm{HH})}$ and $p$-wave scattering volume $v^{(\mathrm{HH})}$ between the heavy atoms, obtained from Fig. 1 and Eqs. (8) (13) (14). The numerical data with different values of $R_{\mathrm{min}}$ collapse into universal curves. Therefore, the three-body parameter of the Efimov states are universally characterized by ($a^{(\mathrm{HH})}$, $r_{\mathrm{vdw}}$) for the bosons and ($v_{p}^{(\mathrm{HH})}$, $r_{\mathrm{vdw}}$) for the fermions, respectively. In other words, they are insensitive to the short-range details, but are universally characterized by the two-body scattering parameters which describe the long-range asymptotic behavior of the two heavy atoms. The numerical data in Fig. 2 excellently agree with the curves of Eqs. (15) (16), vindicating our QDT analytical formulae. Some points in Fig. 2 deviating from the curves are higher angular momentum states, whose energies change rapidly as $R_{\mathrm{min}}$ hence $a^{(\mathrm{HH})}$ and $v^{(\mathrm{HH})}$ are varied; although they significantly deviate from the analytical curves of Eqs. (15) (16), the data points with different $R_{\mathrm{min}}$ still tend to collapse into a single curve (Fig. 2 (b)(c)) thanks to the angular-insensitive nature of the short-range phase originating from $R_{\mathrm{min}}\ll r_{\mathrm{vdw}}$ [52]. The marginal disagreement of the large binding energy region $|E|\simeq 1/Mr_{\mathrm{vdw}}^{2}$ in Fig. 2 (c) is ascribed to the breakdown of the low-energy condition used to derive Eqs. (15) (16).
@@ III.2 Weak dipole interaction $a_{dd}\ll r_{\mathrm{vdw}}$
As shown in Fig. 3, the numerical results generally get shifted toward larger binding energy. This can be understood by the perturbation theory: we have analytically found that the energy shift induced by the dipole interaction is $E-E^{(0)}\propto a_{dd}^{2}$ for the bosons predominantly in the $L=0$ channel, and $E-E^{(0)}\propto a_{dd}$ for the fermions predominantly in the $L=1$ channel, with both negative shifts $E-E^{(0)}<0$ (see Appendix. C). The difference in the scaling originates from the absence (presence) of the diagonal matrix element of the dipole interaction for $L=0$ ($L=1$) states: $\langle L=0|V_{dd}|L=0\rangle=0$, $\langle L=1|V_{dd}|L=1\rangle\neq 0$. This is confirmed in Fig. 4: the change in binding energy, i.e. three-body parameter of the Efimov states, scales quadratically for the bosons, and linearly for the fermions. Notably, in most previous studies on the Efimov effect, the role of the statistics of the particles has been rather marginal, at best to modify the value of $s_{\ell}$, and thereby change the absence/presence of the Efimov effect through the condition $s_{\ell}^{2}<0$ as the mass ratio is varied [2, 35, 49, 50, 3, 4, 5, 6]. In other words, the statistics of the particles has been essentially irrelevant once its effect is incorporated into the value of $s_{\ell}$. Figure 4 is one of the few examples where the quantum statistics of the particles plays an explicit role in qualitatively changing the universal behavior of the Efimov states.
@@ III.3 Moderate dipole interaction $a_{dd}\sim r_{\mathrm{vdw}}$
Figure 4: Energy shift induced via the dipole interaction with a fixed value of $R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.40$ for (a) the bosons (blue circles) and (b) fermions (orange squares). The dashed lines (black) are the perturbation results in Eqs (51) (53), where the unperturbed energy $E^{(0)}$ and the matrix elements of Eq. (50) are evaluated with the energies and wavefunctions numerically obtained for the purely van der Waals system in Sec. III.1.
@@
Figure 5: Binding energy of the bosonic Efimov states for the mass ratio $M/m=27.5855...$ corresponding to 166Er-166Er-6Li system. Weaker dipole strength $a_{dd}/r_{\mathrm{vdw}}=0.40$ (top row (a)(b)(g)(h)), realistic dipole strength for 166Er-166Er of $a_{dd}/r_{\mathrm{vdw}}=0.86755....$ (middle row (c)(d)(i)(j)), and stronger dipole strength $a_{dd}/r_{\mathrm{vdw}}=1.5$ (bottom row (e)(f)(k)(l)). (a)-(f) The horizontal axis is the $s$-wave scattering length between the heavy atoms estimated by Eq. (8), i.e. without the dipole interaction. (g)-(l) The horizontal axis is the $s$-wave scattering length between the heavy atoms numerically obtained by solving the coupled-channel two-body equation in Eq. (17). The left (right) panels in each column show the negative (positive) $s$-wave scattering length. The solid curve (black) is the analytical QDT formula for non-dipolar system in Eq. (15).
@@
Figure 6: $S$-wave scattering length between the heavy atoms as $R_{\mathrm{min}}$ is varied, obtained numerically by solving Eq. (17). The solid curve (black) is the $s$-wave scattering length without the dipole interaction $a^{(\mathrm{HH})}_{\mathrm{no-dd}}$, i.e., the analytical formula of Eqs (13) (8). The dotted vertical lines (black) are the same as those in Fig. 1(a).
@@
In Fig. 5(g)-(l), we show the binding energies of the Efimov states plotted against $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. Similar to Fig. 5(a)-(f), the data with different values of $R_{\mathrm{min}}$ mostly collapse into a single curve, demonstrating the universality. One notable difference is that the agreement between the numerical data and the purely van der Waals analytical curves (black) is much better in (g)-(l) than those in (a)-(f): except for the broad avoided crossing at $a_{\mathrm{with-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$, a significant fraction of the numerical results agree well with the van der Waals curves not only for the weak dipole interaction, but also for the realistic dipole strength (Fig. 5(i)(j)) and stronger dipole interaction ((k)(l)). This suggests that the van der Waals quantum QDT of the bosonic Efimov states in Eq. (15) can capture the universal behavior of the Efimov states in the presence of the dipole interaction, once the effects of the dipole interaction are incorporated into the $s$-wave scattering length between the heavy atoms: $a^{(\mathrm{HH})}=a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. In other words, the major role of the dipole interaction is to change, hence renormalize the value of the $s$-wave scattering length between the heavy atoms, so that the van der Waals universality of the Efimov states in Eq. (15) persists with the renormalized $s$-wave scattering length $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. This renormalized van der Waals universality for the dipolar bosonic Efimov states we have found here seems non-pertubative; the dipole interaction is so large that it significantly changes both the binding energies (Fig. 5(c)-(f)) and the $s$-wave scattering length (Fig. 6(b)(c))) from the purely van der Waals system. Some horizontal data points in Fig. 5(g)(l) which are not present in (a)-(f) are artefacts of the higher-partial-wave induced resonances: the two-body problem between the heavy atoms in Eq. (17) not only shows the $s$-wave dominated resonances, but also shows the resonant enhancement of $a^{(\mathrm{HH})}$ induced by higher-partial-wave channels [59, 61], whose approximate positions are indicated by the vertical black dotted lines in Fig. 6. The narrow nature of these higher-partial-wave induced resonances leads to an abrupt change of $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$ as $R_{\mathrm{min}}$ is changed, resulting in flat energy spectra in Fig. 5(g)-(l). Equation (15) cannot explain these flat data points. We note however that these flat energy spectra are absent in Fig. 5(a)-(b), and they should therefore be interpreted as artefacts originaing from the two-body effect of $a^{(\mathrm{HH})}$ rather than a genuine three-body phenomenon. Physically, when we attempt to evaluate the binding energies of the Efimov states from an experimentally obtained value of $a^{(\mathrm{HH})}$ (see Sec. IV), we can simply use the van der Waals curves of Eq. (15) by neglecting the stray flat spectra, except for the case where the two Er atoms are coincidentally in the proximity of the $s$-wave resonance induced by a $d$-wave or higher-angular-momentum molecular state. In the avoided crossing region, the variance of the spectrum is larger than in the other regions (see Fig. 5(l)). This is due to the finite values of $R_{\mathrm{min}}$, as it can be checked that reducing it further does lead to a converged universal behavior. For the finite values of $R_{\mathrm{min}}$ used in the figures (which is varied between 0.25 and 0.4 $r_{\mathrm{vdw}}$), it is still justified to call this region’s behavior universal despite the noticeable energy variance because it remains much smaller than the Efimov period, in sheer contrast with the strength of the van der Waals interaction at short distance which varies by a factor 16. Yet, a $d$-wave dimer of the two heavy atoms appearing around $a_{\mathrm{with-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$ (see Fig. 11 in Appendix. D where the dimer energies are presented) seems to significantly affect two-body and three-body behaviors in this region, making the simple renormalized van der Waals universality of Eq. (15) inadequate.
@@
While the physical origin of the renormalized van der Waals universality is yet to be clarified, we suspect that the angular momentum insensitive nature of the short-range phase [52] is playing a crucial role; we have confirmed in our three-body calculations that a substantial fraction of the wavefunction is occupying $\ell\geq 2$ states. If the short-range phase shifts in various angular momentum channels are directly related with each other, as in the angular-momentun-insensitive QDT [52], $a^{(\mathrm{HH})}$ characterizing the $s$-wave short-range phase can universally capture the scatterings occuring in all the angular momentum channels. We also suspect that a universal three-body repulsion appearing at moderately large distance may be another possible mechanism [23, 21, 22, 46, 47]; in the three-body system of identical bosons interacting via the van der Waals interaction, the appearance of a non-adiabatic three-body repulsion at a large distance $r\simeq 2r_{\mathrm{vdw}}$ is ascribed to be the physical origin of the universality of the three-body parameter of the Efimov states [23, 21, 22]. Furthermore, for three identical bosons interacting via purely dipolar interactions, a three-body repulsion has also been found to appear universally at $r\simeq a_{dd}$ [46]. The multi-channel effects with an interplay of the van der Waals and dipole interaction in our sysetm may lead to a similar three-body repulsion scenario. We also remark that our universality of the Efimov states involving higher partial-wave channels is analogous to the partial-wave phase-locking mechanism found in the resonant exchange collisions [62, 63], where the phase shifts of different scattering channels between the atoms are related with each other.
@@
Figure 7: $K_{c,F}^{(3D)}$ between two heavy fermionic atoms for $a_{dd}/r_{\mathrm{vdw}}=0.86755....$ corresponding to the realistic dipole strength between 167Er atoms ($a_{dd}=75.7a_{0}$ and $r_{\mathrm{vdw}}=65.9a_{0}$ [57, 58]). The solid curve (black) is $K_{c,F}^{(3D)}$ calculated with purely dipole interaction (i.e. without the van der Waals interaction) in Eq. (18): $K_{c,F}^{(3D)}=J_{3}\left(q_{\mathrm{min}}\right)/Y_{3}\left(q_{\mathrm{min}}\right)$ with $q_{\mathrm{min}}=\sqrt{\frac{48a_{dd}}{5R_{\mathrm{min}}}}$
@@
where $J_{3}$ and $Y_{3}$ are the Bessel and Neuman functions corresponding to two independent solutions under the central $1/r^{3}$ potential. $K_{c,F}^{(3D)}$ is introduced in an analogous manner as the QDT parameter $K^{c}$ in the van der Waals case (see Eq. (5)), and hence is expected to be a universal parameter characterizing the short-range two-body phase shift. In Fig. 7, we show $K_{c,F}^{(3D)}$ numerically calculated from Eqs. (18) (19) with the hard-wall boundary condition $u_{1}(r=R_{\mathrm{min}})=0$. We note that the van der Waals term is essential in determining the short-range phase: $K_{c,F}^{(3D)}$ is markedly different from that obtained by neglecting the van der Waals term (black solid), suggesting both the van der Waals and dipole interactions are relevant for characterizing the scattering between the two heavy fermionic atoms.
@@
Figure 8: Binding energy of the fermionic Efimov states for the mass ratio $M/m=27.7520...$ corresponding to 167Er-167Er-6Li system. $K_{c,F}^{(3D)}$ in the horizontal axis is obtained numerically from Eqs. (18) (19). Realistic dipole strength for 167Er-167Er of $a_{dd}/r_{\mathrm{vdw}}=0.8705....$ (top row (a)(b)), and stronger dipole strengths of $a_{dd}/r_{\mathrm{vdw}}=1.6$ (middle row (c)(d)), and $a_{dd}/r_{\mathrm{vdw}}=2.0$ (bottom row (e)(f)) are presented.

### v3

@@ III Numerical Results
We show the numerical solutions of Eq. (1) in this section. As we are particularly interested in Er-Er-Li three-body systems, we perform the calculations with Er-Li mass ratio: $M/m=27.5855...$ for the bosonic system corresponding to 166Er-6Li, and $M/m=27.7520...$ for the fermions corresponding to 167Er-6Li, respectively. The differential equation (1) is discretized in the log-scaled coordinate space between $R_{\mathrm{min}}$ and $R_{\mathrm{max}}$, where $R_{\mathrm{max}}$ is typically taken as $\sim$ 400-1000 $r_{\mathrm{vdw}}$. We find that 4000 grid points and $\ell_{\mathrm{max}}=$10-20 are enough to attain convergence for most cases, while we take up to 30000 grid points and $\ell_{\mathrm{max}}\approx 40$ for some cases of stronger dipole interactions.
@@ III.1 Non-dipolar case $a_{dd}=0$
Figure 2: Binding energy of the Efimov states plotted against (a)(b) the $s$-wave scattering length $a^{(\mathrm{HH})}$ for the bosons, and (c)(d) $p$-wave scattering volume $v_{p}^{(\mathrm{HH})}$ for the fermions, obtained by converting $R_{\mathrm{min}}$ in Fig. 1 into $a^{(\mathrm{HH})}$ and $v_{p}^{(\mathrm{HH})}$ using Eqs. (8) (13), and (14). The left (right) pannels are negative (positive) scattering length/volume sides, which should smoothly connect with each other in the unitary limit $01/a^{(\mathrm{HH})}=$ 0, $01/v_{p}^{(\mathrm{HH})}=$ 0 corresponding to the border between the left and right panels. The solid curves (black) are the analytical results of Eqs. (15) and (16), and the dashed curves (green) are the dimer binding energy of the heavy atoms. The vertical dotted lines (black) are the same as those in Fig. 1.
@@
First, we show the results without the dipole interaction: in Fig. 1, we show the binding energy of the Efimov states (in the van der Waals energy unit $1/Mr_{\mathrm{vdw}}^{2}$) as the short-range boundary $R_{\mathrm{min}}$ is varied. As $R_{\mathrm{min}}$ gets smaller, the short-range van der Waals attraction increases, leading to the successive appearance of bound states and increase of the binding energy. The numerical data points almost perfectly agree with the analytical curves obtained with the QDT in Eq. (7), particularly at small energy. There are a few points which significantly deviate from the analytical curves. They are ascribed to higher angular momentum states of $\ell=2,4,...$ for the bosons and $\ell=3,5,...$ for the fermions, respectively. The thresholds at which the $\ell=2$ ($\ell=3$) bound states start to appear, denoted by the vertical dotted lines in Fig. 1 (a) and Fig. 1 (b), agree with most of the stray data points, supporting this interpretation. A few stray points which agree with neither are ascribed to higher angular momentum channels $\ell\geq 4$ ($\ell\geq 5$). This classification of the states has also been confirmed from the shape of the wave functions. Because the higher angular momentum states are localized to shorter distances than the $s$-wave ($p$-wave) state owing to the centrifugal barrier, they are more sensitive to $R_{\mathrm{min}}$, resulting in the steeper binding energy curves. The binding energy curves of different partial waves do not render an avoided crossing but simply cross because the angular momentum is a good quantum number in the absence of the dipole interaction.
@@
In Fig. 2, we show the binding energies plotted against the $s$-wave scattering length $a^{(\mathrm{HH})}$ and $p$-wave scattering volume $v^{(\mathrm{HH})}$ between the heavy atoms, obtained from Fig. 1 and Eqs. (8) (13) (14). The numerical data with different values of $R_{\mathrm{min}}$ collapse into universal curves. Therefore, the three-body parameter of the Efimov states is universally characterized by ($a^{(\mathrm{HH})}$, $r_{\mathrm{vdw}}$) for the bosons and ($v_{p}^{(\mathrm{HH})}$, $r_{\mathrm{vdw}}$) for the fermions, respectively. In other words, they are insensitive to the short-range details, but are universally characterized by the two-body scattering parameters which describe the long-range asymptotic behavior of the two heavy atoms. The numerical data in Fig. 2 excellently agree with the curves of Eqs. (15) and (16), vindicating our QDT analytical formulas. Some points in Fig. 2 deviating from the curves are higher angular momentum states, whose energies change rapidly as $R_{\mathrm{min}}$ hence $a^{(\mathrm{HH})}$ and $v_{p}^{(\mathrm{HH})}$ are varied; although they significantly deviate from the analytical curves of Eqs. (15) and (16), the data points with different $R_{\mathrm{min}}$ still tend to collapse into a single curve (Fig. 2 (b)(c)) due to the angular-momentum-insensitive nature of the short-range phase originating from $R_{\mathrm{min}}\ll r_{\mathrm{vdw}}$ [52]. The marginal disagreement of the large binding energy region $|E|\simeq 1/Mr_{\mathrm{vdw}}^{2}$ in Fig. 2 (c) is ascribed to the breakdown of the low-energy condition used to derive Eqs. (15) and (16).
@@ III.2 Weak dipole interaction $a_{dd}\ll r_{\mathrm{vdw}}$
As shown in Fig. 3, the numerical results generally get shifted toward larger binding energy. This can be understood by the perturbation theory: we have analytically found that the energy shift induced by the dipole interaction is $E-E^{(0)}\propto a_{dd}^{2}$ for the bosons predominantly in the $L=0$ channel, and $E-E^{(0)}\propto a_{dd}$ for the fermions predominantly in the $L=1$ channel, with both negative shifts $E-E^{(0)}<0$ (see Appendix. C). The difference in the scaling originates from the absence (presence) of the diagonal matrix element of the dipole interaction for $L=0$ ($L=1$) states: $\langle L=0|V_{dd}|L=0\rangle=0$, $\langle L=1|V_{dd}|L=1\rangle\neq 0$. This is confirmed in Fig. 4: the change in binding energy, i.e., three-body parameter of the Efimov states, scales quadratically for the bosons, and linearly for the fermions. Notably, in most previous studies on the Efimov effect, the role of the statistics of the particles has been rather marginal, at best to modify the value of $s_{\ell}$, and thereby change the absence/presence of the Efimov effect through the condition $s_{\ell}^{2}<0$ as the mass ratio is varied [2, 35, 49, 50, 3, 4, 5, 6]. In other words, the statistics of the particles has been essentially irrelevant once its effect is incorporated into the value of $s_{\ell}$. Figure 4 is one of the few examples where the quantum statistics of the particles plays an explicit role in qualitatively changing the universal behavior of the Efimov states.
@@ III.3 Moderate dipole interaction $a_{dd}\sim r_{\mathrm{vdw}}$
Figure 4: Energy shift induced via the dipole interaction with a fixed value of $R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.40$ for (a) the bosons (blue circles) and (b) fermions (orange squares). The dashed lines (black) are the perturbation results in Eqs. (51) and (53), where the unperturbed energy $E^{(0)}$ and the matrix elements of Eq. (50) are evaluated with the energies and wave functions numerically obtained for the purely van der Waals system in Sec. III.1.
@@
Figure 5: Binding energy of the bosonic Efimov states for the mass ratio $M/m=27.5855...$ corresponding to 166Er-166Er-6Li system. Weaker dipole strength $a_{dd}/r_{\mathrm{vdw}}=0.40$ (top row (a)(b)(g)(h)), realistic dipole strength for 166Er-166Er of $a_{dd}/r_{\mathrm{vdw}}=0.86755....$ (middle row (c)(d)(i)(j)), and stronger dipole strength $a_{dd}/r_{\mathrm{vdw}}=1.5$ (bottom row (e)(f)(k)(l)). (a)-(f) The horizontal axis is the $s$-wave scattering length between the heavy atoms estimated by Eq. (8), i.e., without the dipole interaction. (g)-(l) The horizontal axis is the $s$-wave scattering length between the heavy atoms numerically obtained by solving the coupled-channel two-body equation in Eq. (17). The left (right) panels in each column show the negative (positive) $s$-wave scattering length. The solid curve (black) is the analytical QDT formula for non-dipolar system in Eq. (15).
@@
Figure 6: $S$-wave scattering length between the heavy atoms as $R_{\mathrm{min}}$ is varied, obtained numerically by solving Eq. (17). The solid curve (black) is the $s$-wave scattering length without the dipole interaction $a^{(\mathrm{HH})}_{\mathrm{no-dd}}$, i.e., the analytical formula of Eqs. (13) and (8). The dotted vertical lines (black) are the same as those in Fig. 1(a).
@@
In Fig. 5(g)-(l), we show the binding energies of the Efimov states plotted against $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. Similar to Fig. 5(a)-(f), the data with different values of $R_{\mathrm{min}}$ mostly collapse into a single curve, demonstrating the universality. One notable difference is that the agreement between the numerical data and the purely van der Waals analytical curves (black) is much better in (g)-(l) than those in (a)-(f): except for the broad avoided crossing at $a_{\mathrm{with-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$, a significant fraction of the numerical results agree well with the van der Waals curves not only for the weak dipole interaction, but also for the realistic dipole strength (Fig. 5(i)(j)) and stronger dipole interaction ((k)(l)). This suggests that the van der Waals QDT of the bosonic Efimov states in Eq. (15) can capture the universal behavior of the Efimov states in the presence of the dipole interaction, once the effects of the dipole interaction are incorporated into the $s$-wave scattering length between the heavy atoms: $a^{(\mathrm{HH})}=a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. In other words, the major role of the dipole interaction is to change, hence renormalize, the value of the $s$-wave scattering length between the heavy atoms, so that the van der Waals universality of the Efimov states in Eq. (15) persists with the renormalized $s$-wave scattering length $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$. This renormalized van der Waals universality for the dipolar bosonic Efimov states we have found here seems non-pertubative; the dipole interaction is so large that it significantly changes both the binding energies (Fig. 5(c)-(f)) and the $s$-wave scattering length (Fig. 6(b)(c))) from the purely van der Waals system. Some horizontal data points in Fig. 5(g)(l) which are not present in (a)-(f) are artifacts of the higher-partial-wave-induced resonances: the two-body problem between the heavy atoms in Eq. (17) not only shows the $s$-wave dominated resonances, but also shows the resonant enhancement of $a^{(\mathrm{HH})}$ induced by higher-partial-wave channels [59, 61], whose approximate positions are indicated by the vertical black dotted lines in Fig. 6. The narrow nature of these higher-partial-wave-induced resonances leads to an abrupt change of $a^{(\mathrm{HH})}_{\mathrm{with-dd}}$ as $R_{\mathrm{min}}$ is changed, resulting in flat energy spectra in Fig. 5(g)-(l). Equation (15) cannot explain these flat data points. We note, however, that these flat energy spectra are absent in Fig. 5(a)-(b), and they should therefore be interpreted as artifacts originating from the two-body effect of $a^{(\mathrm{HH})}$ rather than a genuine three-body phenomenon. Physically, when we attempt to evaluate the binding energies of the Efimov states from an experimentally obtained value of $a^{(\mathrm{HH})}$ (see Sec. IV), we can simply use the van der Waals curves of Eq. (15) by neglecting the stray flat spectra, except for the case where the two Er atoms are coincidentally in the proximity of the $s$-wave resonance induced by a $d$-wave or higher-angular-momentum molecular state. In the avoided crossing region, the variance of the spectrum is larger than in the other regions (see Fig. 5(l)). This is due to the finite values of $R_{\mathrm{min}}$, as it can be checked that reducing it further does lead to a converged universal behavior. For the finite values of $R_{\mathrm{min}}$ used in the figures (which is varied between 0.25 and 0.4 $r_{\mathrm{vdw}}$), it is still justified to call this region’s behavior universal despite the noticeable energy variance because it remains much smaller than the Efimov period, in sheer contrast with the strength of the van der Waals interaction at short distance, which varies by a factor 16. Yet, a $d$-wave dimer of the two heavy atoms appearing around $a_{\mathrm{with-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$ (see Fig. 11 in Appendix. D where the dimer energies are presented) seems to significantly affect two-body and three-body behaviors in this region, making the simple renormalized van der Waals universality of Eq. (15) inadequate.
@@
While the physical origin of the renormalized van der Waals universality is yet to be clarified, we suspect that the angular-momentum-insensitive nature of the short-range phase [52] is playing a crucial role; we have confirmed in our three-body calculations that a substantial fraction of the wave function is occupying $\ell\geq 2$ states. If the short-range phase shifts in various angular momentum channels are directly related with each other, as in the angular-momentun-insensitive QDT [52], $a^{(\mathrm{HH})}$ characterizing the $s$-wave short-range phase can universally capture the scatterings occuring in all the angular momentum channels. We also suspect that a universal three-body repulsion appearing at moderately large distance may be another possible mechanism [23, 21, 22, 46, 47]; in the three-body system of identical bosons interacting via the van der Waals interaction, the appearance of a non-adiabatic three-body repulsion at a large distance $r\simeq 2r_{\mathrm{vdw}}$ is ascribed to be the physical origin of the universality of the three-body parameter of the Efimov states [23, 21, 22]. Furthermore, for three identical bosons interacting via purely dipolar interactions, a three-body repulsion has also been found to appear universally at $r\simeq a_{dd}$ [46]. The multi-channel effects with an interplay of the van der Waals and dipole interaction in our system may lead to a similar three-body repulsion scenario. We also remark that our universality of the Efimov states involving higher partial-wave channels is analogous to the partial-wave phase-locking mechanism found in the resonant exchange collisions [62, 63], where the phase shifts of different scattering channels between the atoms are related with each other.
@@
Figure 7: $K_{c,F}^{(3D)}$ between two heavy fermionic atoms for $a_{dd}/r_{\mathrm{vdw}}=0.86755....$ corresponding to the realistic dipole strength between 167Er atoms ($a_{dd}=75.7a_{0}$ and $r_{\mathrm{vdw}}=65.9a_{0}$ [57, 58]). The solid curve (black) is $K_{c,F}^{(3D)}$ calculated with purely dipole interaction (i.e., without the van der Waals interaction) in Eq. (18): $K_{c,F}^{(3D)}=J_{3}\left(q_{\mathrm{min}}\right)/Y_{3}\left(q_{\mathrm{min}}\right)$ with $q_{\mathrm{min}}=\sqrt{\frac{48a_{dd}}{5R_{\mathrm{min}}}}$.
@@
where $J_{3}$ and $Y_{3}$ are the Bessel and Neumann functions corresponding to two independent solutions under the central $1/r^{3}$ potential. $K_{c,F}^{(3D)}$ is introduced in an analogous manner as the QDT parameter $K^{c}$ in the van der Waals case (see Eq. (5)), and hence is expected to be a universal parameter characterizing the short-range two-body phase shift. In Fig. 7, we show $K_{c,F}^{(3D)}$ numerically calculated from Eqs. (18) and (19) with the hard-wall boundary condition $u_{1}(r=R_{\mathrm{min}})=0$. We note that the van der Waals term is essential in determining the short-range phase: $K_{c,F}^{(3D)}$ is markedly different from that obtained by neglecting the van der Waals term (black solid), suggesting both the van der Waals and dipole interactions are relevant for characterizing the scattering between the two heavy fermionic atoms.
@@
Figure 8: Binding energy of the fermionic Efimov states for the mass ratio $M/m=27.7520...$ corresponding to 167Er-167Er-6Li system. $K_{c,F}^{(3D)}$ in the horizontal axis is obtained numerically from Eqs. (18) and (19). Realistic dipole strength for 167Er-167Er of $a_{dd}/r_{\mathrm{vdw}}=0.8705....$ (top row (a)(b)), and stronger dipole strengths of $a_{dd}/r_{\mathrm{vdw}}=1.6$ (middle row (c)(d)), and $a_{dd}/r_{\mathrm{vdw}}=2.0$ (bottom row (e)(f)) are presented.